OpenStudy (anonymous):

Please help! (u(v)^2)^3 / v^2u^3

OpenStudy (campbell_st):

simplify the numerator 1st it uses the power of a power rule \[(x^a)^b = x^{a \times b}\] so the numerator is \[u^{1 \times 3}v^{2 \times 3} \] simplify this, then use the index law for division \[\frac{x^a}{x^b} = x^{a - b}\] hope it helps

OpenStudy (anonymous):

thank you!!